OpenStudy (anonymous):

Find the magnitude and direction angle of r=-48i+6j. I think I understand how to get the magnitude but not understanding the direction angle...

OpenStudy (anonymous):

So to get the magnitude would be \[\sqrt{-48^{2}+6^{2}}\]

OpenStudy (anonymous):

I think.

OpenStudy (anonymous):



OpenStudy (anonymous):

you are right on the magnitude

OpenStudy (anonymous):

So the direction angle goes back to a unit circle?

OpenStudy (anonymous):

Or am I just crazy

OpenStudy (anonymous):

yeah it pretty much does, all the unit circle is is a way to look at angles, so take tan(x) = opp/ adj tan(x) = 6/48 <-- 48 is positive because it's just a magnitude the angle in the triangle is atan(6/48)

OpenStudy (anonymous):

so the direction from positive x is 180 - that angle found

OpenStudy (anonymous):

Okay then would the magnitude be kept in a fraction form 'cause the decimal goes on forever o.o

OpenStudy (anonymous):



OpenStudy (anonymous):

the magnitude is up to you or however your teacher wants it. it is safe to round off usually to 2 decimals

OpenStudy (anonymous):

:) okay awesome that makes sense! Thanks a lot.

OpenStudy (anonymous):

no problem, glad to help